OpenStudy (anonymous):

solve: 5 over 6 = x over 24

OpenStudy (anonymous):

I am not so familiar with English mathematic expressions as it is not my basic language. If you will upload a paint picture of the expression written, I will do my best to help.

OpenStudy (anonymous):

5/6 = x/24 you cross multiply 5(24) = 6x 120=6x divide 6 to eliminate on the other side 120/6=6x/6 x=20